Consider the following sample data:
x 32 23 18 36 13 40 35 24
y 31 37 34 60 25 45 43 38

Click here for the Excel Data File

b.

Calculate b1 and b0. What is the sample regression equation? (Round intermediate calculations to at least 4 decimal places and final answers to 2 decimal places.)

  yˆy^  +  x
c.

Find the predicted value for y if x equals 11, 16, and 21. (Round intermediate coefficient values and final answers to 2 decimal places.)

yˆy^       
  If x = 11    
  If x = 16    
  If x = 21      
Answer #1

Given clata can be represented in table form. T x2 32 37 34 60 1600 24 Х ху 31 :992 1024 23 85 l 529 18 61 2 324 36 2160 1296we use we want to predict y when X=11,16821. sample regression equation. ģ= 1.61+ 1,358 % 0 put x=11 in eano. ŷ = 1.61 + (1.3